WHAT IS THE x-INTERCEPT AND y-INTERCEPT OF THE EQUATION 5x-3y=30?
Found 2 solutions by jai_kos, uma:
Answer by jai_kos(139)

5x-3y=30--->(1)
The x intercept is given by putting y =0 in the equation(1), we get
5x =30
x =30/5
x =6
The x intercept is 6
The y intercept is given by putting x =0 in the equation(1), we get
-3y = 30
y =30 /-3
y = -10

Therefoer the "x" intercept is 6 and the y intercept is -10

Answer by uma(370)

The given equation is 5x - 3y = 30
To find the x intercept replace y by 0.
==> 5x = 30
==> 5x/5 = 30/5
==> x = 6
To find the y intercept replace x by 0.
==> 0 - 3y = 30
==> -3y = 30
==> -3y/-3 = 30/-3
==> y = - 10
Thus the x intercept = 6 and y intercept = - 10
Good Luck!!!